# Doped resonating valence bond states: How robust are the spin ice phases in 3D Rydberg arrays

### v1

@@ I Introduction
Figure 2: Pyrochlore lattice. (a) Each sites in the pyrochlore lattice is shared by an up-pointing tetrahedron and a down-pointing tetrahedron. Different colored dots indicate that the atoms in ground states or Rydberg state. The configuration satisfies the ice rule. (b) The green dashed line connects the centers of the nearest-neighbour tetrahedron to form a hexagonal cell. We label the atoms at different positions from 1 to 6 respectively. The atoms in the Rydberg state are mapped to dimers, while the atoms in the ground state are mapped to empty links. Therefore, the Rydberg atomic array on the pyrochlore lattice can be mapped to a QDM on diamond lattice. (c) The pyrochlore lattice is a compound lattice composed of four face-centered cubic (FCC) lattices. Each vertex of a tetrahedron belongs to a different sublattice. The sites in different sublattices are labeled as 1, 2, 3, and 4. (d) For one sublattice, the direct lattice vectors are $\overrightarrow{a_{1}},\overrightarrow{a_{2}},\overrightarrow{a_{3}}$. (e) The reciprocal lattice is body-centered cubic (BCC) with vectors are $\overrightarrow{b_{1}},\overrightarrow{b_{2}},\overrightarrow{b_{3}}$. We measure the dispersion along the $\overrightarrow{b_{1}}+\overrightarrow{b_{3}}$ and $\overrightarrow{b_{3}}$.
@@
$\langle O\rangle=\frac{\langle G|O|G\rangle}{\langle G|G\rangle}=\frac{\sum_{C,C^{\prime}}\langle C|O|C^{\prime}\rangle}{\sum_{C,C^{\prime}}\langle C|C^{\prime}\rangle},$ (6)
$\frac{\langle G|O|G\rangle}{\langle G|G\rangle}=\frac{\sum_{C}\langle C|O|C^{\prime}\rangle}{\sum_{C}\langle C|C\rangle},$ (7)
here we only samples $C$ because the $C^{\prime}$ is fixed once the $O$ and $C$ are known. Therefore, the expectation value of observables could be gained via sampling different configurations $C$ with equal weight. That is, randomly sample the configurations $C$ and average the values $\langle C|O|C^{\prime}\rangle$ to gain $\langle O\rangle$.
Through the above way, the static properties can be calculated and explored. For the dynamical excitations above the RK ground state, the single mode approximation (SMA) is a useful tool for analyzing the low-lying dispersion of the model [65, 66, 67, 68, 69, 70, 71, 72, 73], which transforms the problems of solving dynamical properties into evaluations of equal-time correlations as can be achieved in Eq. (7). For example, for the dynamical dimer density correlation, we assume that the excited-states are described by the single-mode ansatz: $|\textbf{q},\alpha\rangle=D_{\alpha}(\textbf{q})|G\rangle$, then then the upper bound of the excited state energy $\omega_{SMA}$ can be obtained as:
where $D_{\alpha}(\mathbf{q})=\frac{1}{\sqrt{N}}\sum_{\mathbf{r}}e^{-i\mathbf{q}\cdot\mathbf{r}}D_{\alpha}(\mathbf{r})$ is the Fourier transformed dimer density operator in momentum space, and $D_{\alpha}(\mathbf{r})$ is the dimer density operator in the real space. Here each dimer is labeled by its unit cell position with the sublattice index $\alpha$ and postion r, respectively. We only consider the dispersion in a same sublattice here.
Figure 3: Dispersion of dimer density correlation for the original RK-QDM obtained by SMA for the diamond lattice ($24\times 24\times 24$) with strict one-dimer-per-site (star) and two-dimers-per-site (circle) along the Brillouin zone path $\Gamma(0,0,0)\rightarrow(0,4\pi,0)$ and $\Gamma\rightarrow(2\pi,2\pi,-2\pi)$.
Based on SMA, we consider a $24\times 24\times 24$ diamond lattice with perfect one-dimer-per-site and two-dimers-per-site, and evaluate the dimer dispersion along the direction of $\overrightarrow{b_{1}}+\overrightarrow{b_{3}}$ and $\overrightarrow{b_{3}}$ in momentum space based on Eq. (8). The computational approach is to sample the dimer configurations with equal-weight through a random walk. All the updates are local loops which cannot change the winding sector of the dimer configurations. This process is repeated $10000$ times, and the results are statistically averaged to approximate the dispersion of the pyrochlore lattice, $\omega_{SMA}$. We calculate the dispersion of each of the four sublattices separately, and then average them to obtain the dispersion of the entire system. The result is plotted in Fig. 3. The quantum excitation at the RK point exhibits only one gapless points located at $\Gamma(0,0,0)$ (also equivalent to $(0,4\pi,0)$ and $(2\pi,2\pi,-2\pi)$). This only gapless mode corresponds to the photon excitations as predicted by theory. Since the RK point is described by the Lifshitz’s theory with dynamical exponent $z=2$, its dispersion is quadratic [74, 39].
If the configurations are not perfectly carrying fixed integer numbers of dimers per site, then the doping case can be discussed. For example, when we add dimers into the system of one-dimer-per-site and set the soft constraint with one or two dimers per site only, the site with two dimers can be treated as a doped point. The density of doping is defined as the percentage of the number of doped sites that make up the ground state relative to the total number of lattice sites. In the following, we will tune the local filling from one-dimer-per-site to two, even to three. Now we have to define the monomers under doping. If a site violates the local dimer constraint of the reference, it is a monomer. In the paper, when doping density is less than $50\%$, the rule for choosing the reference filling is: one-dimer-per-site if the local dimer filling is one or two; two dimers per site if the local filling is two or three; three dimers per site if the local filling is three or four. When the doping density is greater than $50\%$, we choose the upper local filling as the reference. For example, for one-dimer-per-site configuration with doping density $>50\%$, we choose the reference filling is two-dimers-per-site, the site with one dimer becomes a monomer.
In the gauge field language, monomers are regarded as magnetic charge fluctuations respect to the background reference configurations. If we incorporate quantum dynamics of monomers as their hopping, we can construct a RK-like effective Hamiltonian based on the original RK-QDM, supplemented by the dimer hopping term and the related diagonal term:
[…]
where the $t^{\prime}$ and $V^{\prime}$ terms run over all the nearest-neighbour links of the lattice. The third term is the kinetic energy term for exchanging dimers between adjacent links with the coupling strength $-t^{\prime}$, and the fourth term is the corresponding potential energy term with the coupling strength $V^{\prime}$. This system exhibits a global $U(1)$ symmetry associated with conservation of total dimers:
@@
Figure 4: The dispersion of dimer density operator of the doped QDM at the RK point $t=t^{\prime}=V=V^{\prime}=1$. We take the Brillouin zone path $\Gamma\leavevmode\nobreak\ (0,0,0)\rightarrow(0,4\pi,0)$ and $\Gamma\rightarrow(2\pi,2\pi,-2\pi)$ on a diamond lattice with system size $24\times 24\times 24$. Different colors correspond to (a) one-dimer-per-site with varying doping densities ($0.904\%$, $25\%$, $49.913\%$, $70.023\%$ and $95.833\%$) and (b) two-dimers-per-site with varying doping densities ($0.904\%$, $25\%$, and $49.913\%$) .
@@
We further investigate the dynamical dimer-dimer correlation of the doped QDM Eq. (9) under the one-dimer-per-site with varying doping densities ($0.904\%,25\%,49.913\%,70.023\%$ and $95.833\%$) and two-dimers-per-site with varying doping densities ($0.904\%,25\%$ and $49.913\%$), see Fig. 4. The results indicate that doping of monomers gaps out the photon excitations at the $\Gamma$ point by Higgsing the emergent $U(1)$ gauge field. Consequently, the Coulomb liquid phase vanishes and the system should become trivially confined upon doping.
@@
Figure 5: Off-diagonal dimer correlation of the doped QDM at the RK point $t=t^{\prime}=V=V^{\prime}=1$. The circles represent the off-diagonal correlations along the $\overrightarrow{a_{1}}$ direction as a function of the distance $|i-j|$, Different colors correspond to (a) one-dimer-per-site with the varying doping density ($0\%,4.167\%,25\%,49.913\%$ and $83.333\%$) and (b) two-dimers-per-site with the varying doping density ($0\%,4.167\%,25\%,$ and $49.913\%$) , where $0\%$ corresponds to applying a strong constraint to the system that prohibits the presence of monomers. Here, the distance between nearest-neighbour links in the sublattice is set to 1.
@@
To further explore the physical properties of the QDM upon doping, we evaluate the off-diagonal dimer correlations of the doped QDM by taking $O=D_{i}^{+}D_{j}^{-}$ in Eq. (7). This dynamical correlation function measures excitations involving creation or annihilation of dimers in the RK ground state. Here, we set the distance between nearest-neighbour links in a same sublattice to 1.
@@
In Fig. 5, we present the off-diagonal correlations along the $\overrightarrow{a_{1}}$ direction in one sublattice, under the background of one-dimer-per site and two-dimers-per-site, and compare the impact of different doping densities ($4.167\%,25\%,49.913\%,83.333\%$ and $0\%$) on the off-diagonal correlation. The results show that the off-diagonal correlation vanishes in the undoped case, as the dimer hopping term $D_{i}^{+}D_{j}^{-}$ breaks up the one-dimer-per-site local constraint. In contrast, while such local constraints become soft upon doping, the system immediately develops non-vanishing dimer off-diagonal long-range correlations. This presence of off-diagonal long-range order can be understood as condensation of monomers that disrupts the local constraint and higgses the $U(1)$ gauge field, hence the Coulomb liquid breaks down.
@@
Figure 6: The vector directions associated with the dual (diamond) lattice. (a) The Rydberg array located on the sites of pyrochlore lattice can be mapped to a dimer on a link of dual (diamond) lattice which is composed of two face-centered cubic lattices. The sites of the dual (diamond) lattice are the body centers of each upward and downward tetrahedron, and they belong to different sublattices. We label the different sublattices as 1 and 2. (b) For one sublattice, the direct lattice vectors are $\overrightarrow{a}_{1}^{{}^{\prime}},\overrightarrow{a}_{2}^{{}^{\prime}},\overrightarrow{a}_{3}^{{}^{\prime}}$. (c) The reciprocal lattice is body-centered cubic (BCC) with vectors are $\overrightarrow{b}_{1}^{{}^{\prime}},\overrightarrow{b}_{2}^{{}^{\prime}},\overrightarrow{b}_{3}^{{}^{\prime}}$.
@@
Figure 7: Dispersion of monomer density correlation at the RK point ($t=t^{\prime}=V=V^{\prime}=1$) along the Brillouin zone path $\Gamma\rightarrow(0,4\pi,0)$ and $\Gamma\rightarrow(2\pi,2\pi,-2\pi)$ on a $24\times 24\times 24$ diamond lattice. Different colors correspond to (a) one-dimer-per-site with doping densities ($0.90422\%,4.167\%,45.8333\%,54.1667\%$ and $70.023\%$) and (b) two-dimers-per-site with doping densities ($0.90422\%,4.167\%$ and $45.8333\%$).
@@
where $m(\mathbf{q})=\frac{1}{\sqrt{N}}\sum_{\mathsf{r}}e^{-i\mathbf{q}\cdot\mathsf{r}}m_{\alpha}(\mathsf{r})$ represents the density of monomer in momentum space, and $\frac{1}{\sqrt{N}}$ is the normalization constant. $m_{\alpha}(\mathsf{r})$ is the density of monomer in real space, where the label $\alpha$ refers to a specific sublattice of the composite lattice. $m_{\alpha}(\mathsf{r})=1$ means that the site at position $\mathsf{r}$ in a given configuration $C$ violates the background constraint, while $m_{\alpha}(\mathsf{r})=0$ indicates that the site at position $\mathsf{r}$ in configuration $c$ follows the background constraint. Here, we also consider a $24\times 24\times 24$ dual diamond lattice with local constraint based on the Eq. (11).
In Fig. 7, we present the dispersions of monomers based on one-dimer-per-site configurations with varying doping density ($0.90422\%,4.167\%,45.8333\%,54.1667\%$ and $70.023\%$) and two-dimers-per-site configurations with varying doping density ($0.90422\%,4.167\%$ and $45.8333\%$). The results reveal gapless states at the $\Gamma$ point. This gapless mode is identified as the Goldstone mode associated with the monomer condensates that also order at the $\Gamma$ point. This phase is quite analogous to superfluids as it also exhibits a gapless Goldstone mode associated with $U(1)$ symmetry breaking. However, it is still different from superfluids, as the Goldstone mode here exhibits a quadratic dispersion with vanishing superfluid stiffness, and is unstable against viscosity.

### v2

@@ I Introduction
Figure 2: Pyrochlore lattice. (a) Each sites in the pyrochlore lattice is shared by an up-pointing tetrahedron and a down-pointing tetrahedron. Different colored dots indicate that the atoms are in ground states or Rydberg states. The configuration satisfies the ice rule. (b) The pink dashed line connects the centers of the nearest-neighbor tetrahedron to form a hexagonal cell. We label the atoms at different positions from 1 to 6 respectively. The atoms in the Rydberg state are mapped to dimers, while the atoms in the ground state are mapped to empty links. Therefore, the Rydberg atomic array on the pyrochlore lattice can be mapped to a QDM on diamond lattice. (c) The pyrochlore lattice is a composite lattice composed of four face-centered cubic (FCC) lattices. Each vertex of a tetrahedron belongs to a different sublattice. The sites in different sublattices are labeled as 1, 2, 3, and 4. (d) For one sublattice, the direct lattice vectors are $\overrightarrow{a_{1}},\overrightarrow{a_{2}},\overrightarrow{a_{3}}$. (e) The reciprocal lattice is body-centered cubic (BCC) with vectors $\overrightarrow{b_{1}},\overrightarrow{b_{2}},\overrightarrow{b_{3}}$. We measure the dispersion along the $\overrightarrow{b_{1}}+\overrightarrow{b_{3}}$ and $\overrightarrow{b_{3}}$.
Because the ground state wave-function at the RK point is an equal-weight superposition of all possible configurations, i.e., $\mbox{$|G\rangle$}=\sum_{C}\mbox{$|C\rangle$}$, the expectation value of any physical observable $O$ can be sampled as
$\langle O\rangle=\frac{\langle G|O|G\rangle}{\langle G|G\rangle}=\frac{\sum_{C,C^{\prime}}\langle C|O|C^{\prime}\rangle}{\sum_{C,C^{\prime}}\langle C|C^{\prime}\rangle}.$ (6)
$\frac{\langle G|O|G\rangle}{\langle G|G\rangle}=\frac{\sum_{C}\langle C|O|C^{\prime}\rangle}{\sum_{C}\langle C|C\rangle}.$ (7)
In this case, we only sample $C$, because once $C$ is fixed, there is only one $C^{\prime}$ such that the matrix element $\langle C^{\prime}|O|C\rangle\neq 0$. Therefore, the expectation value of observables could be gained via sampling different configurations $C$ with equal weight. That is, randomly sample the configurations $C$ and average the values $\langle C|O|C^{\prime}\rangle$ to gain $\langle O\rangle$.
Through the above way, the static properties can be calculated and explored. For the dynamical excitations above the RK ground state, the single mode approximation (SMA) is a useful tool for analyzing the low-lying dispersion of the model [70, 71, 72, 73, 74, 75, 76, 77, 78], which transforms the problems of solving dynamical properties into evaluations of equal-time correlations, as shown in Eq. (7). For example, for the dynamical dimer density correlation, we assume that the excited-states are described by the single-mode ansatz: $|\textbf{q},\alpha\rangle=D_{\alpha}(\textbf{q})|G\rangle$, then then the upper bound of the excited state energy $\omega_{SMA}$ can be obtained as:
where $D_{\alpha}(\mathbf{q})=\frac{1}{\sqrt{N}}\sum_{\mathbf{r}}e^{-i\mathbf{q}\cdot\mathbf{r}}D_{\alpha}(\mathbf{r})$ is the Fourier transformed dimer density operator in momentum space, and $D_{\alpha}(\mathbf{r})$ is the dimer density operator in the real space. Here, each dimer is labeled by its unit cell position with the sublattice index $\alpha$ and position r, respectively. We only consider the dispersion in the same sublattice here.
Figure 3: Dispersion of dimer density correlation for the original RK-QDM obtained by SMA for the diamond lattice ($24\times 24\times 24$) with strict one dimer per site (star) and two dimers per site (circle) along the Brillouin zone path $\Gamma(0,0,0)\rightarrow(0,4\pi,0)$ and $\Gamma\rightarrow(2\pi,2\pi,-2\pi)$.
Based on SMA, we consider a $24\times 24\times 24$ diamond lattice with perfect one dimer per site and two dimers per site, and evaluate the dimer dispersion along the direction of $\overrightarrow{b_{1}}+\overrightarrow{b_{3}}$ and $\overrightarrow{b_{3}}$ in momentum space based on Eq. (8). The computational approach is to sample the dimer configurations with equal-weight through a random walk. All the updates are local loops which cannot change the winding sector of the dimer configurations. This process is repeated $10000$ times, and the results are statistically averaged to approximate the dispersion of the pyrochlore lattice, $\omega_{SMA}$. We calculate the dispersion of each of the four sublattices separately, and then average them to obtain the dispersion of the entire system. The result is plotted in Fig. 3. The quantum excitation at the RK point exhibits only one gapless points located at $\Gamma(0,0,0)$ (also equivalent to $(0,4\pi,0)$ and $(2\pi,2\pi,-2\pi)$). This only gapless mode corresponds to the photon excitations as predicted by theory. Since the RK point is described by the Lifshitz’s theory with dynamical exponent $z=2$, its dispersion is quadratic [79, 40].
If the configurations do not perfectly carry fixed integer numbers of dimers per site, then the doping case can be discussed. For example, when we add dimers into the system of one dimer per site and set the soft constraint with one or two dimers per site only, the site with two dimers can be treated as a doped point. The density of doping is defined as the percentage of the number of doped sites that make up the ground state relative to the total number of lattice sites. In the following, we will tune the local filling from one-dimer-per-site to two, even to three. Now we have to define the monomers under doping. If a site violates the local dimer constraint of the reference, it is a monomer. In the paper, when doping density is less than $50\%$, the rule for choosing the reference filling is: one dimer per site if the local dimer filling is one or two; two dimers per site if the local filling is two or three; three dimers per site if the local filling is three or four. When the doping density is greater than $50\%$, we choose the upper local filling as the reference. For example, for one-dimer-per-site configuration with doping density $50\%$, we choose the reference filling as two dimers per site, the site with one dimer becomes a monomer.
In the gauge field language, monomers are regarded as magnetic charge fluctuations with respect to the background reference configurations. If we incorporate the quantum dynamics of monomers as their hopping, we can construct an RK-like effective Hamiltonian based on the original RK-QDM, supplemented by the dimer hopping term and the related diagonal term:
@@
where the $t^{\prime}$ and $V^{\prime}$ terms run over all the nearest-neighbor links of the lattice. The third term is the kinetic energy term for exchanging dimers between adjacent links with the coupling strength $t^{\prime}$, and the fourth term is the corresponding potential energy term with the coupling strength $V^{\prime}$. This system exhibits a global U(1) symmetry associated with conservation of total dimers:
@@
Figure 4: The dispersion of dimer density operator of the doped QDM at the RK point $t=t^{\prime}=V=V^{\prime}=1$. We take the Brillouin zone path $\Gamma\penalty 10000\ (0,0,0)\rightarrow(0,4\pi,0)$ and $\Gamma\rightarrow(2\pi,2\pi,-2\pi)$ on a diamond lattice with system size $24\times 24\times 24$. Different colors correspond to (a) one dimer per site with varying doping densities ($0.904\%$, $25.0\%$, $49.9\%$, $70.0\%$ and $95.8\%$) and (b) two dimers per site with varying doping densities ($0.904\%$, $25.0\%$, and $49.9\%$) .
@@
We further investigate the dynamical dimer-dimer correlation of the doped QDM Eq.(9) under the one dimer per site with varying doping densities ($0.904\%,25.0\%,49.9\%,70.0\%$ and $95.8\%$) and two dimers per site with varying doping densities ($0.904\%,25.0\%$ and $49.9\%$), see Fig. 4. The results indicate that doping of monomers gaps out the photon excitations at the $\Gamma$ point by Higgsing the emergent U(1) gauge field. Consequently, the Coulomb liquid phase vanishes and the system should become trivially confined upon doping.
@@
Figure 5: Off-diagonal dimer correlation of the doped QDM at the RK point $t=t^{\prime}=V=V^{\prime}=1$. The circles represent the off-diagonal correlations along the $\overrightarrow{a_{1}}$ direction as a function of the distance |ij||i-j|, Different colors correspond to (a) one dimer per site with the varying doping density ($0\%,4.17\%,25.0\%,49.9\%$ and $83.3\%$) and (b) two dimers per site with the varying doping density ($0\%,4.17\%,25.0\%,$ and $49.9\%$) , where $0\%$ corresponds to applying a strong constraint to the system that prohibits the presence of monomers. Here, the distance between nearest-neighbor links in the sublattice is set to 1. The values of the off-diagonal dimer correlation are significantly larger than their corresponding errors, the error bars are obscured by the circular data markers.
@@
To further explore the physical properties of the QDM upon doping, we evaluate the off-diagonal dimer correlations of the doped QDM by taking $O=D_{i}^{+}D_{j}^{-}$ in Eq. (7). This dynamical correlation function measures excitations involving creation or annihilation of dimers in the RK ground state. Here, we set the distance between nearest-neighbor links in the same sublattice to 1.
In Fig. 5, we present the off-diagonal correlations along the $\overrightarrow{a_{1}}$ direction in one sublattice, under the background of one-dimer-per site and two-dimers-per-site, and compare the impact of different doping densities ($4.17\%,25.0\%,49.9\%,83.3\%$ and $0\%$) on the off-diagonal correlation. We did not include the point at |ij|=0|i-j|=0 ($\langle D_{i}^{+}D_{j}^{-}\rangle=1$) in the figure. This omission is intentional, as the steep drop from ij=0i-j=0 to ij=1i-j=1 would cause all the curves for $i-j\geq 1$ to collapse onto each other, which would obscure the differences between curves for different doping densities.
In addition, when $i-j\geq 1$, the value of the off-diagonal dimer correlation $\langle D_{i}^{+}D_{j}^{-}\rangle$ is independent of the distance |ij||i-j|. The underlying reason for this phenomenon is that the distribution of monomers within the same sublattice is statistically uncorrelated. According to Eqs. (6) and (7), only dimer configurations $C$ with site $i$ not attached to any monomers and site $j$ attached to two monomers has non-vanishing contribution to $\langle D_{i}^{+}D_{j}^{-}\rangle$. Therefore, $\langle D_{i}^{+}D_{j}^{-}\rangle$ essentially measures the statistical weight of such configurations. Given that the wave function at the generalized RK point is the equal-weight superposition of all allowed dimer configurations, As a result, monomers are spatially independently distributed, leading to a correlation function that is independent of |ij||i-j|, and hence a flat curve.
@@
The results show that the off-diagonal correlation vanishes in the undoped case, as the dimer hopping term $D_{i}^{+}D_{j}^{-}$ breaks up the one-dimer-per-site local constraint. In contrast, while such local constraints become soft upon doping, the system immediately develops non-vanishing dimer off-diagonal long-range correlations. This presence of off-diagonal long-range order can be understood as condensation of monomers that disrupts the local constraint and breaks the U(1) gauge field, hence the Coulomb liquid breaks down.
@@
Figure 6: The vector directions associated with the dual (diamond) lattice. (a) The Rydberg array located on the sites of the pyrochlore lattice can be mapped to a dimer on a link of the dual (diamond) lattice which is composed of two face-centered cubic lattices. The sites of the dual (diamond) lattice are the body centers of each upward and downward tetrahedron, and they belong to different sublattices. We label the different sublattices as 1 and 2. (b) For one sublattice, the direct lattice vectors are $\overrightarrow{a}_{1}^{{}^{\prime}},\overrightarrow{a}_{2}^{{}^{\prime}},\overrightarrow{a}_{3}^{{}^{\prime}}$. (c) The reciprocal lattice is body-centered cubic (BCC) with vectors are $\overrightarrow{b}_{1}^{{}^{\prime}},\overrightarrow{b}_{2}^{{}^{\prime}},\overrightarrow{b}_{3}^{{}^{\prime}}$.
@@
Figure 7: Dispersion of monomer density correlation at the RK point ($t=t^{\prime}=V=V^{\prime}=1$) along the Brillouin zone path $\Gamma\rightarrow(0,4\pi,0)$ and $\Gamma\rightarrow(2\pi,2\pi,-2\pi)$ on a $24\times 24\times 24$ diamond lattice. Different colors correspond to (a) one dimer per site with doping densities ($0.904\%,4.17\%,45.8\%,54.2\%$ and $70.0\%$) and (b) two dimers per site with doping densities ($0.904\%,4.17\%$ and $45.8\%$).
@@
where $m(\mathbf{q})=\frac{1}{\sqrt{N}}\sum_{\mathsf{r}}e^{-i\mathbf{q}\cdot\mathsf{r}}m_{\alpha}(\mathsf{r})$ represents the density of monomer in momentum space, and $\frac{1}{\sqrt{N}}$ is the normalization constant. $m_{\alpha}(\mathsf{r})$ is the density of monomer in real space, where the label $\alpha$ refers to a specific sublattice of the composite lattice (Fig. 6). $m_{\alpha}(\mathsf{r})=1$ means that the site at position $\mathsf{r}$ in a given configuration $C$ violates the background constraint, while $m_{\alpha}(\mathsf{r})=0$ indicates that the site at position $\mathsf{r}$ in configuration $c$ follows the background constraint. Here, we also consider a $24\times 24\times 24$ dual diamond lattice with local constraint based on the Eq. (11).
In Fig. 7, we present the dispersions of monomers based on one-dimer-per-site configurations with varying doping density ($0.904\%,4.17\%,45.8\%,54.2\%$ and $70.0\%$) and two-dimers-per-site configurations with varying doping density ($0.904\%,4.17\%$ and $45.8\%$). The results reveal gapless states at the $\Gamma$ point. This gapless mode is identified as the Goldstone mode associated with the monomer condensates that also order at the $\Gamma$ point. This phase is quite analogous to superfluids as it also exhibits a gapless Goldstone mode associated with U(1) symmetry breaking. However, it is still different from superfluids, as the Goldstone mode here exhibits a quadratic dispersion with vanishing superfluid stiffness, and is unstable against viscosity.